OpenStudy (anonymous):

which of the following is equivalent to 6^(3/2)

OpenStudy (anonymous):

sqrt(2)^9

OpenStudy (anonymous):

sqrt(6)^3

OpenStudy (anonymous):

3sqrt(6)^2

OpenStudy (anonymous):

3^3

Nnesha (nnesha):

exponent rule \[\huge\rm \sqrt[n]{x^m } = x^\frac{ m }{ n }\]

OpenStudy (anonymous):

6^(3/2) = 6^(1 + 1/2) = 6*6^(1/2)

OpenStudy (anonymous):

3sqrt(6)^2

Nnesha (nnesha):

nope focus! :-)